OpenStudy (anonymous):

At Southside Middle School, 40% of the students play an instrument. There are 305 students. How many students play an instrument?

OpenStudy (anonymous):

305 x . 40 = number of students that play an instrument

OpenStudy (godlovesme):

% means /100 so 45/100 * 305 .45* 305..? Got it

OpenStudy (some.random.cool.kid):

@above is correct

OpenStudy (anonymous):

@bittybitt is correct

OpenStudy (godlovesme):

@smartkid2000

OpenStudy (triciaal):

set up a ratio to find equivalent percent means per 100 let the number of students that play = x 45/100 = x/305 solve for x as above x = (45/100)*305